Proposal:
K Kids
Target of Oportunity:
No
Time Observed:
12.25
Time Engineering:
0.0
Time lost technical:
0.0
Time lost weather:
0.0
Time lost other:
0.0
Primary Investigator:
Todd Henry
Email:
Institution:
SMARTS
Instrument:
CTIO 1.5m Echelle
Organization:
SMARTS
Total Program Time:
12.25

At the very Beginning of the Night: ThAr Fiber lamps 1116 and 1117 few counts: used PI_LAMPS.vi and then got the right lamp, file 1118.
File 1166 ThAr Slicer no counts : PI_LAMPS.vi shows switch UP but led OFF, similar to what happened to 1159 (switch UP led OFF) but then the lamp was OK.
File 1170 'HD209290' Tel err ~ +5 min in DEC, stopped: resent it, ok. N wind ~ 20 MPH. Then noted the PI_LAMPS.vi reported ThAr lamp was still ON so, turned it OFF.
Intermittent DIMM, 06:20 to 07:48 UT.
Extra 15' integration of 'TYC_9126-1532-1' ended with the Sun at -11.6 degrees of altitude.